You deposit $1250 in an account that pays 5.4% annual interest compounded monthly.
a. When does your balance first exceed $1750?
after _____ years and _____ months

Answer by Boreal(15235)

1750=1250(1+0.054/12)^12t, where t=number of years.
1750/1250=1.4=(1.0045)^12t
logs both sides
log(1.4)=12tlog(1.0045)
12t=log(1.4)/log(1.0045)=74.94
It will be in the 75th month or 6y3m.
check 1250(1.0045)^75=$1750.47